OpenStudy (anonymous):

How many solutions does the system of equations have? 2x + 3y = 6 y= -2/3x-1

OpenStudy (anonymous):

you just have to solve for x and y first

OpenStudy (anonymous):

But what numbers do i plug in?

OpenStudy (anonymous):

try plugging in (0 , 0)

OpenStudy (anonymous):

(0 , 0) | | x , y solve for y when x is 0 solve for x when y is 0

OpenStudy (texaschic101):

2x + 3y = 6 y = -2/3x - 1 (sub -2/3x - 1 in for y in the 1st equation) 2x + 3y = 6 2x + 3(-2/3x -1)= 6 2x - 2x - 3 = 6 2x - 2x = 6 + 3 0 = 9 This system of equations has NO SOLUTIONS